OpenStudy (anonymous):

If |u| = 10, |v| = 8, and the angle formed between them is 60°, then u · v = ___?

OpenStudy (jhannybean):

\[\vec u \cdot \vec v = |\vec u||\vec v|\cos(\theta)\]

OpenStudy (jhannybean):

Plug in all your givens and solve

OpenStudy (anonymous):

10*8*0.5

OpenStudy (anonymous):

40?

OpenStudy (anonymous):

yeap

OpenStudy (anonymous):

thank you